OpenStudy (anonymous):

Magnitude of force acting on spring with spring constant of 275 N/m that is stretched 14.3cm? As always, any help is great(:

OpenStudy (anonymous):

do you know the expression for spring force?

OpenStudy (08surya):

\[F=(275\times14.3)\div100\]

OpenStudy (08surya):

AND answer will be in newton

OpenStudy (theeric):

I'll just explain how @08surya arrived at that conclusion. I agree, but also there is the Code of Conduct ( http://openstudy.com/code-of-conduct) where we agree to explain solutions, also. The formula for spring force, as Mashy asked of you, is \(F=-k\ x\). \(F\) is the force, \(k\) is the spring constant, \(x\) is the distance stretched or compressed (I'll use stretched), and the negative sign says that the force will be in the opposite direction of the stretching. You are given \(k=275\ [N/m]\), and \(x=14.3\ [cm]\). But, for the units to work out, we want those centimeters to be meters. \(14.3\ [cm]\times\dfrac{1\ [m]}{100\ [cm]}=14.3\ [m]\div100\) seems to be what 08surya did. That's just a unit conversion, leaving the numbers separate but canceling the units. You can just go straight to \(14.3\ [m]\div 100=.143\ [m]\), but I won't, because I want to get to 08surya's solution. So you plug \(k\) and \(x\) into the \(F=-k\ x\) formula to get \(F=-275\ [N/m]\times 14.3\ [m]\div100\) We can put in parenthesis to show that \(F=-\left(275\ [N/m]\times 14.3\ [m]\right)\div100\) Now, units cancel. I'll just throw the units at the end. \(F=-\Large\left(\normalsize\left(275\times 14.3\right)\div100\right)\normalsize\ [N]\) The only difference now is the negative sign. 08surya decided to find the magnitude of the spring force. I used a formula that gives direction. Since I considered the distance of stretching to be positive, and the force is negative, the force is in the direction of compression. It is trying to compress the spring back to where it was. If you want to throw big words around, you would say that spring force is a \(\sf restoring\ force\). :)

OpenStudy (08surya):

@theeric-buddy u r right i just solved the question

OpenStudy (anonymous):

Thanks to both! I really appreciate it! (: